# Robust Neural Information Retrieval: An Adversarial and Out-of-distribution Perspective

### v1

@@ Abstract.
Recent advances in neural information retrieval (IR) models have significantly enhanced their effectiveness over various IR tasks. The robustness of these models, essential for ensuring their reliability in practice, has also garnered significant attention. With a wide array of research on robust IR being proposed, we believe it is the opportune moment to consolidate the current status, glean insights from existing methodologies, and lay the groundwork for future development. We view the robustness of IR to be a multifaceted concept, emphasizing its necessity against adversarial attacks, out-of-distribution (OOD) scenarios and performance variance. With a focus on adversarial and OOD robustness, we dissect robustness solutions for dense retrieval models (DRMs) and neural ranking models (NRMs), respectively, recognizing them as pivotal components of the neural IR pipeline. We provide an in-depth discussion of existing methods, datasets, and evaluation metrics, shedding light on challenges and future directions in the era of large language models. To the best of our knowledge, this is the first comprehensive survey on the robustness of neural IR models, and we will also be giving our first tutorial presentation at SIGIR 2024  [^1] The tutorial slides about robust IR at SIGIR 2024 are available at https://sigir2024-robust-information-retrieval.github.io/.. Along with the organization of existing work, we introduce a Benchmark for robust IR (BestIR), a heterogeneous evaluation benchmark for robust neural information retrieval, which is publicly available at https://github.com/Davion-Liu/BestIR. We hope that this study provides useful clues for future research on the robustness of IR models and helps to develop trustworthy search engines  [^2] A curated list of researches related to robust IR can be found at https://github.com/Davion-Liu/Awesome-Robustness-in-Information-Retrieval..
Robustness, Information Retrieval, Survey
@@ 1. Introduction
Information retrieval (IR) plays a pivotal role in numerous real-world applications, such as web search (Ilan, 1998), digital libraries (Chowdhury and Chowdhury, 2003) and e-commerce search (Karmaker Santu et al., 2017). According to the global overview report from Digital 2023, nearly 82% of Internet users between 18 and 64 have used a search engine or web portal in the past month.https://datareportal.com/reports/digital-2023-global-overview-report Specifically, IR is the process of finding and providing relevant information in response to the user query from a large collection of data. Recently, with advances in deep learning, neural IR models have witnessed significant progress (Guo et al., 2020, 2022). With the development of training methodologies such as pre-training (Gao and Callan, 2022; Ma et al., 2022) and fine-tuning (Qu et al., 2021; Zhan et al., 2021; Khattab and Zaharia, 2020), neural IR models have demonstrated remarkable effectiveness in learning query-document relevance patterns.
@@
In real-world deployment of neural IR models, an aspect equally essential as their effectiveness is their robustness. A good IR system must not only exhibit high effectiveness under normal conditions but also demonstrate robustness in the face of abnormal conditions. The natural openness of IR systems makes them vulnerable to intrusion, and the consequences can be severe. For example: (i) Search engines are vulnerable to black hat SEO attacks,https://www.bleepingcomputer.com/news/security/15-000-sites-hacked-for-massive-google-seo-poisoning-campaign/ necessitating significant efforts to curb these infringements.https://www.bbc.com/news/technology-28687513, https://developers.google.com/search/docs/essentials (ii) In addition, search engines are confronted with large amounts of unseen data on a daily basis. The working algorithm needs to be updated frequently to ensure that search effectiveness is maintained.https://developers.google.com/search/news
Figure 1. The statistics of publications related to robust neural information retrieval.
@@
Recently, academics have begun to investigate the robustness of IR systems (Wu et al., 2022b; Thakur et al., 2021; Cohen et al., 2018; Liu et al., 2022a, 2023d; Chen et al., 2023c). As neural networks gain increasing popularity in IR, many studies have found that IR systems that are based on neural networks inherit a wide variety of robustness issues from deep neural network. The field of robust neural IR is garnering increasing attention, as evidenced by the growing number of papers published annually, as depicted in Figure 1. The robustness issues are differently represented in real IR scenarios and raise concerns about deploying neural IR systems into the real world. Therefore, the study of robust neural IR is crucial for building reliable IR systems.
What is the definition of robustness in IR?
@@
In this paper, we begin by defining robustness in the context of IR. In IR, it is well-accepted that, user attention mainly focuses on the Top-$K$ results and increases with higher rankings (Niu et al., 2012). Based on this, we argue that robustness in IR refers to the consistent performance and resilience on Top-$K$ results of an IR system when faced with a variety of unexpected scenarios. Robustness is not a simple concept; it encompasses multiple dimensions, as highlighted by research within the machine learning (ML) community (Shafique et al., 2020; Zhang et al., 2020). Specifically, in the field of IR, we identify several facets of robustness: (i) Performance varianceemphasizes the worst-case performance across different individual queries under the independent and identically distributed (IID) data assumption (Wu et al., 2022b); (ii) Out-of-distribution (OOD) robustnessrefers to the generalizability of an IR model on unseen queries and documents from different distributions of the training dataset (Thakur et al., 2021); and (iii) Adversarial robustnessrefers to the ability of the IR model to defend against malicious adversarial attacks aimed at manipulating rankings (Wu et al., 2022b).
@@
In this survey, our focus is on adversarial robustness and OOD robustness, which have garnered significant attention. (i) For adversarial robustness, studies primarily approach the topic from two angles, i.e., adversarial attack and defense, to enhance the robustness of IR models. (ii) For OOD robustness, the emphasis is on improving the generalizability of IR models to both unseen documents and unseen queries. Additionally, we zero in on two key components of the neural IR framework: first-stage retrieval and the subsequent ranking stage. More specifically, we focus on dense retrieval models (DRMs) and neural ranking models (NRMs) to further explore the aforementioned research perspectives.
The relation to other surveys
@@
There are a number of surveys on robustness in the field of natural language processing (NLP) (Wang et al., 2022c, 2023) and computer vision (CV) (Drenkow et al., 2021; Akhtar et al., 2021). However, the field of IR presents its own unique characteristics: (i) Unlike NLP, which often focuses on individual examples, IR involves ranking a collection of documents, highlighting the need for robustness across the entire list of rankings. (ii) Differing from continuous image data in CV, IR deals with robustness concerning discrete text documents. Consequently, the studies explored in these surveys are not directly transferrable as references within the IR field.
@@
Surveys specific to the IR domain tend to concentrate on areas like pre-training (Zhao et al., 2024; Fan et al., 2022), ranking models (Guo et al., 2020), initial retrieval stages (Guo et al., 2022), and the explainability of IR systems (Anand et al., 2023b). Yet, there’s a noticeable gap in the literature: a dedicated survey that consolidates and introduces research pertaining to robustness in IR is markedly absent.
@@
On top of the work discussed above, this survey adds the following: (i) Our work complements robustness surveys from other fields, taking into account the distinctive characteristics of different domains. (ii) Our survey complements IR surveys by not only prioritizing effectiveness but also emphasizing robustness. Our ultimate aim is to contribute to help advance reliable and efficient IR systems.
@@
  • A comprehensive overview and categorization: We define robustness in IR by summarizing existing literature and further dividing it into distinct categories.
  • A detailed discussion of methodologies and datasets: We offer a detailed discussion of methodologies, datasets, and evaluation metrics pertinent to each aspect of robustness. Moreover, we integrate the existing datasets mentioned in this survey and propose the BestIR benchmark to facilitate subsequent work.
  • Identification of open issues and future trends: We highlight challenges and potential future trends, particularly in the age of large language models (LLMs).
@@
Figure 2 depicts the organization of our survey: In Section 2, we introduce the IR task, and give a definition and taxonomy of robustness in IR. In Section 3, we examine adversarial attack and defense tasks, alongside their respective datasets, evaluation criteria, and cutting-edge methodologies. In Section 4, we show two key scenarios for OOD robustness, i.e., OOD generalizability on unseen documents and OOD generalizability to unseen queries, and present specific datasets, evaluation metrics, and methods for solving these scenarios. In Section 5, we describe remaining challenges and emerging opportunities for robustness of IR in the era of LLMs. Finally, in Section 6, we summarize the survey and offer concluding remarks.
Figure 2. Overview of the survey. Section 5 only is partially listed here because of space limitations.
@@ 2. Definition and Taxonomy
We use $f$ to denote the IR model, which predicts the relevance score $f(q,d)$ based on a given query $q$ and document $d$. The IR model $f$ is derived by learning from the following objective
@@
To achieve the IR goal, we employ the ranking function $f$ to generate a score for any query-document pair $(q,d)$, reflecting the relevance degree between $q$ and $d$. This allows us to produce a rank $\pi_{f}\left(q,d\right)$ for each $d\in D$ according to the predicted score. In general, the ranking function $f$ can be further abstracted by the following unified formulation
@@
where $\psi$ and $\phi$ are representation functions that extract features from $q$ and $d$ respectively, $\eta$ is the interaction function that extracts features from the pair $(q,d)$, and $g$ is an evaluation function that computes the relevance score based on the feature representations. For neural IR models, in most cases, the matching functions $\psi$, $\phi$, $\eta$, and $g$ are encoded in the network structure so that all can be learned from training data (Guo et al., 2020).
The ranking performance of the IR model is usually evaluated by a metric $M$ focus on the top-$K$ ranking results, e.g., Recall@$K$, NDCG@$K$ and MRR@$K$. Given a triple $\left(q,D,Y\right)$ and an IR model $f$, the score of $M$ on query $q$ is calculated by
@@
Further, given a test dataset $\mathcal{D}_{\mathrm{test}}$, the ranking performance $\mathcal{R}_{M}$ against metric $M$ is calculated by
@@
IR models. In order to balance efficiency and effectiveness, the IR task is usually addressed as a pipelined consisting of a first-stage retrieval stage and re-ranking stage (Chen et al., 2017). (i) The first-stage retrieval identifies a small set of candidate documents from millions of documents. Therefore, in the first-stage retrieval, $D_{i}$ refers to the entire corpus. Considering the efficiency, the neural IR model, represented by the dense IR model (DRM) (Guo et al., 2022), usually adopts a dual-encoder architecture. In this architecture, the interaction function $\eta$ is often null. (ii) The re-ranking stage generates the final ranked list for a query and a small set of candidate documents (Guo et al., 2020), referred to as $D_{i}$ in this stage. To this end, the neural ranking model (NRM) with cross-encoder architecture is often modeled jointly by all the matching functions.
@@ 2.1. Definition of Robustness in IR
Robustness refers to the ability to withstand disturbances or external factors that may cause a system to malfunction or provide inaccurate results (Huber, 1981). It is important for practical applications, especially in safety-critical scenarios, e.g., medical retrieval (Averbuch et al., 2004), financial retrieval (Hering, 2017), and private information retrieval (Chor et al., 1998). If, for any reason, the IR system behaves abnormally, the service provider can lose time, manpower, opportunities and even credibility. With the development of deep learning, robustness has received much attention in the fields of computer vision (CV) (Bhojanapalli et al., 2021) and natural language processing (NLP) (Wang et al., 2023). The concerns about model robustness in these fields are mainly focused on the test phase. In this scenario, the model is trained on an unperturbed dataset but tested for its performance when exposed to adversarial examples or OOD data (Goodfellow et al., 2015a; Quionero-Candela et al., 2009).
@@
In IR, it is generally accepted that users focus the vast majority of their attention on the top-$K$ search results, as evidenced by a significant drop in traffic and click-through rates further down the list (Niu et al., 2012; Wu et al., 2022a). Moreover, ranking metrics like MRR (Craswell, 2009) and NDCG (Järvelin and Kekäläinen, 2002), which primarily evaluate the effectiveness of these top-ranking results. The relationship between Top-K result stability and robust IR is shown in Figure 3. Taking SEO as an example, it aims to get its document displayed in the Top-$K$-ranked results. And a robust neural IR model protects its Top-K results from being affected. Consequently, ensuring the integrity and robustness of the top-$K$-ranked results is crucial for deploying IR models into practical web search applications. Based on this, we present a formal definition of top-$K$ robustness in IR by incorporating the original test dataset $\mathcal{D}_{\mathrm{test}}$ from the initial dataset and the unseen test dataset $\mathcal{D}^{*}_{\mathrm{test}}$.
@@
Let $\delta\geq 0$ denote an acceptable error threshold. Given an IR model $f_{\mathcal{D}_{\mathrm{train}}}$ trained on training dataset $\mathcal{D}_{\mathrm{train}}$ with a corresponding testing dataset $\mathcal{D}_{\mathrm{test}}$, an unseen test dataset $\mathcal{D}^{*}_{\mathrm{test}}$, and an acceptable error threshold $\delta$, for the top-$K$ ranking result, if
@@
Note that in Section 2.2, below $\mathcal{D}^{*}_{\mathrm{test}}$ refers to different test datasets depending on the context: $\mathcal{D}^{\prime}\mathrm{test}$ for test dataset with adversarial examples in adversarial robustness, $\tilde{\mathcal{D}}_{\mathrm{test}}$ for test dataset from new domains in OOD robustness.
Figure 4. A taxonomy of the robustness in IR. In this survey, we pay special attention to adversarial robustness and OOD robustness.

### v2

@@ Abstract.
Recent advances in neural information retrieval models have significantly enhanced these models’ effectiveness across various information retrieval tasks. The robustness of these models, which is essential for ensuring their reliability in practice, has also garnered significant attention. With a wide array of research on robust IR being published, we believe it is the opportune moment to consolidate the current status, glean insights from existing methodologies, and lay the groundwork for future development. We view the robustness of information retrieval to be a multifaceted concept, emphasizing its necessity against adversarial attacks, out-of-distribution scenarios, and performance variance. With a focus on adversarial and out-of-distribution robustness, we dissect robustness solutions for dense retrieval models and neural ranking models, respectively, recognizing them as pivotal components of the neural information retrieval pipeline. We provide an in-depth discussion of methods, datasets, and evaluation metrics, shedding light on challenges and future directions in the era of large language models. As side-products of this survey, we release three additional resources: (i) a curated list of publications related to robust IR, [^1] https://github.com/Davion-Liu/Awesome-Robustness-in-Information-Retrieval (ii) a tutorial based on this survey, [^2] https://sigir24-robust-information-retrieval.github.io and (iii) a heterogeneous benchmark for robust information retrieval, BestIR, that collects all known datasets for evaluation information retrieval systems for robustness. [^3] https://github.com/Davion-Liu/BestIR We hope that this study provides useful clues for future research on the robustness of IR models and helps to develop trustworthy search engines.
Robustness, trustworthy systems
@@ 1. Introduction
Recently, with advances in deep learning, neural information retrieval (IR) models have witnessed significant progress (Guo et al., 2020, 2022). With the development of training methodologies such as pre-training (Gao and Callan, 2022; Ma et al., 2022b) and fine-tuning (Qu et al., 2021; Zhan et al., 2021; Khattab and Zaharia, 2020), neural IR models have demonstrated remarkable effectiveness in learning query-document relevance patterns. In deployment of neural IR models, an aspect equally essential as their effectiveness is their robustness. A good IR system must not only exhibit high effectiveness under normal conditions but also demonstrate robustness in the face of abnormal conditions.
@@
The natural openness of IR systems makes them vulnerable to intrusion, and the consequences can be severe. For example: (i) search engines are vulnerable to black hat search engine optimization (SEO) attacks, [^4] https://www.bleepingcomputer.com/news/security/15-000-sites-hacked-for-massive-google-seo-poisoning-campaign/ necessitating significant efforts to curb these infringements. [^5] https://www.bbc.com/news/technology-28687513, https://developers.google.com/search/docs/essentials and (ii) search engines are confronted with large amounts of unseen data on a daily basis. The working algorithm needs to be improved constantly to ensure that search effectiveness is maintained. [^6] https://developers.google.com/search/news
@@
Recently, academic researchers have begun to investigate the robustness of IR systems (Wu et al., 2022b; Thakur et al., 2021; Cohen et al., 2018; Liu et al., 2022a, 2023d; Chen et al., 2023c). As neural networks gain increasing popularity in IR, many studies have found that neural IR systems inherit a wide variety of problematic robustness issues from deep neural networks. In response, the field of robust neural IR is garnering increasing attention, as evidenced by the growing number of papers published annually, as depicted in Figure 1.777See Appendix A for a description of the protocol we followed to select the sources aggregated in Figure 1 and surveyed in this paper. The robustness issues are differently represented in real IR scenarios and raise concerns about deploying neural IR systems into the real world. Therefore, the study of robust neural IR is crucial for building reliable IR systems.
How to defined robustness in IR?
@@
User attention mainly focuses on the Top-$K$ results and achieves with higher rankings (Niu et al., 2012). Based on this, we argue that robustness in IR refers to the consistent performance and resilience on the Top-$K$ results of an IR system when faced with a variety of unexpected scenarios. Robustness is not a simple concept; it encompasses multiple dimensions, as illustrated by research within the machine learning (ML) community (Shafique et al., 2020; Zhang et al., 2020). In IR, we identify several facets of robustness:
  1. (1)
[…]
    IID robustness emphasizes the worst-case performance across different individual queries under the independent and identically distributed (IID) data assumption (Wu et al., 2022b);
  2. (2)
[…]
    OOD robustness refers to the generalizability of an IR model on unseen queries and documents from different distributions of the training dataset (Thakur et al., 2021); and
  3. (3)
    Adversarial robustness refers to the ability of the IR model to defend against malicious adversarial attacks aimed at manipulating rankings (Wu et al., 2022b).
Figure 1. Statistics of publications related to robust neural information retrieval and covered in this survey. “Other” includes arXiv (mostly), TREC, ICDM, NAACL, and TACL.
@@
In this survey, our focus is on adversarial robustness and out-of-distribution (OOD) robustness, which have garnered significant attention: (i) for adversarial robustness, studies primarily approach the topic from two angles, i.e., adversarial attack and defense, to enhance the robustness of IR models, and (ii) for OOD robustness, the emphasis is on improving the generalizability of IR models to both unseen documents and unseen queries. To study the above two aspects, we zoom in on two key components of the neural IR framework: first-stage retrieval and the subsequent ranking stage. We focus on dense retrieval models (DRMs) and neural ranking models (NRMs) to further explore the aforementioned research perspectives.
Relation to other surveys
@@
There are several surveys on robustness in the fields of natural language processing (NLP) (Wang et al., 2022c, 2023) and computer vision (CV) (Drenkow et al., 2021; Akhtar et al., 2021). However, the field of IR presents its own unique characteristics: (i) unlike NLP, which often focuses on individual examples, IR involves ranking a collection of documents, highlighting the need for robustness across the ranked lists, and (ii) different from continuous image data in CV, IR deals with robustness concerning discrete text documents. Consequently, the studies explored in these surveys are not directly transferrable as references within the IR field.
@@
Surveys specific to the IR domain tend to concentrate on areas like pre-training (Zhao et al., 2024; Fan et al., 2022), ranking models (Guo et al., 2020), initial retrieval stages (Guo et al., 2022), and the explainability of IR systems (Anand et al., 2023b). Thus, there is a noticeable gap in the literature: a dedicated survey that consolidates and introduces research pertaining to robustness in IR is absent.
@@
On top of the work discussed above, this survey adds the following: (i) our work complements robustness surveys from other fields, taking into account the distinctive characteristics of different domains, and (ii) our survey complements IR surveys by not only prioritizing effectiveness but also emphasizing robustness. To complement this survey we also release the following resources alongside it: (i) a curated list of publications related to robust IR, (ii) a tutorial on robust IR, and (iii) a benchmark that collects datasets for assessing the robustness of IR systems. See footnotes 1, 2, and 3, respectively.
@@
  • A comprehensive overview and categorization: we define robustness in IR by summarizing the literature and further dividing it into distinct categories; we provide a curated list of publications to support this aspect of the survey.
  • A detailed discussion of methodologies and datasets: we offer a detailed discussion of methodologies, datasets, and evaluation metrics pertinent to each aspect of robustness. We support this part of the survey by providing a tutorial and a benchmark, BestIR, which integrates the datasets mentioned in this survey to facilitate follow-up work.
  • Identification of open issues and future trends: we highlight challenges and potential future trends, particularly in the age of large language models.
@@
Figure 2 depicts the organization of our survey. In Section 2, we introduce the IR task, and give a definition and taxonomy of robustness in IR. In Section 3, we examine adversarial attack and defense tasks, alongside their respective datasets, evaluation criteria, and state-of-the-art methodologies. In Section 4, we show two key scenarios for OOD robustness, i.e., OOD generalizability on unseen documents and OOD generalizability to unseen queries, and present specific datasets, evaluation metrics, and methods for solving these scenarios. In Section 5, we describe remaining challenges and emerging opportunities for robustness of IR in the era of LLMs. Finally, Section 6 summarizes the survey and offers concluding remarks.
Figure 2. Overview of the survey. Section 5 is only partially listed here because of space limitations.
@@ 2. Definition and Taxonomy
We use $f$ to denote the IR model; it predicts the relevance score $f(q,d)$ based on a given query $q$ and document $d$. The IR model $f$ is derived by learning from the following objective:
@@
We employ the ranking function $f$ to generate a score for any query-document pair $(q,d)$, reflecting the relevance degree of $d$ given $q$. This allows us to produce a rank $\pi_{f}\left(q,d\right)$ for each $d\in D$ according to the predicted score. In general, the ranking function $f$ can be further abstracted by the following unified formulation:
@@
where $\psi$ and $\phi$ are representation functions that extract features from $q$ and $d$, respectively, $\eta$ is the interaction function that extracts features from the pair $(q,d)$, and $g$ is an evaluation function that computes the relevance score based on the feature representations. For neural IR models, in most cases, the representation functions $\psi$, $\phi$, $\eta$, and $g$ are encoded in the network structure so that all can be learned from training data (Guo et al., 2020).
The ranking performance of an IR model is usually evaluated by a metric $M$ that focuses on the top-$K$ ranking results, e.g., Recall@$K$, NDCG@$K$ and MRR@$K$. Given a triple $\left(q,D,Y\right)$ and an IR model $f$, the score of $M$ on query $q$ is calculated by:
@@
Further, given a test dataset $\mathcal{D}_{\mathrm{test}}$, the ranking performance $\mathcal{R}_{M}$ against metric $M$ is calculated by:
@@
IR models. In order to balance efficiency and effectiveness, the IR task is usually addressed as a pipeline consisting of a first-stage retrieval stage and re-ranking stage (Chen et al., 2017):
  1. (1)
    The first-stage retrieval identifies a small set of candidate documents from millions of documents. Therefore, in the first-stage retrieval, $D_{i}$ refers to the entire corpus. Considering efficiency, the neural IR model, represented by the dense retrieval model (DRM) (Guo et al., 2022), usually adopts a dual-encoder architecture. In this architecture, the interaction function $\eta$ is often null.
  2. (2)
    The re-ranking stage generates the final ranked list for a query and a small set of candidate documents (Guo et al., 2020), referred to as $D_{i}$ in this stage. To this end, the neural ranking model (NRM) with cross-encoder architecture is often modeled jointly by all the matching functions.
@@ 2.1. Definition of Robustness in IR
Robustness refers to the ability to withstand disturbances or external factors that may cause a system to malfunction or provide inaccurate results (Huber, 1981). It is important for practical applications, especially in safety-critical scenarios, e.g., medical retrieval (Averbuch et al., 2004), financial retrieval (Hering, 2017), and private information retrieval (Chor et al., 1998). If, for any reason, the IR system behaves abnormally, the service provider can lose time, manpower, opportunities and even credibility. With the development of deep learning, robustness has received much attention in the fields of CV (Bhojanapalli et al., 2021) and NLP (Wang et al., 2023). Concerns about model robustness in these fields are mainly focused on the test phase. In this scenario, the model is trained on an unperturbed dataset but tested for its performance when exposed to adversarial examples or OOD data (Goodfellow et al., 2015a; Quionero-Candela et al., 2009).
@@
In most deployed systems, when presented with a ranked list users focus most of their attention on the top-$K$ search results, as evidenced by a significant drop in traffic and click-through rates further down the list (Niu et al., 2012; Wu et al., 2022a) and by the prevalence of ranking metrics like MRR (Craswell, 2009) and NDCG (Järvelin and Kekäläinen, 2002), which primarily evaluate the effectiveness of these top-ranking results. The relationship between Top-$K$ result stability and robust IR is shown in Figure 3. Taking SEO as an example, it aims to get its document displayed in the Top-$K$-ranked results. And a robust neural IR model protects its Top-$K$ results from being affected. Consequently, ensuring the integrity and robustness of the top-$K$-ranked results is crucial for deploying IR models into practical web search applications. Based on this, we present a formal definition of top-$K$ robustness in IR by incorporating the original test dataset $\mathcal{D}_{\mathrm{test}}$ from the initial dataset and the unseen test dataset $\mathcal{D}^{*}_{\mathrm{test}}$.
@@
Let $\delta\geq 0$ denote an acceptable error threshold. Let $f_{\mathcal{D}_{\mathrm{train}}}$ be an IR model trained on training dataset $\mathcal{D}_{\mathrm{train}}$, with a corresponding test dataset $\mathcal{D}_{\mathrm{test}}$, and an unseen test dataset $\mathcal{D}^{*}_{\mathrm{test}}$, for the top-$K$ ranking results. If
@@
In Section 2.2, below, $\mathcal{D}^{*}_{\mathrm{test}}$ refers to different test datasets depending on the context: $\mathcal{D}^{\prime}_{\mathrm{test}}$ for test datasets with adversarial examples in adversarial robustness, $\tilde{\mathcal{D}}_{\mathrm{test}}$ for test dataset from new domains in OOD robustness.
Figure 4. A taxonomy of robustness in IR. In this survey, we pay special attention to adversarial robustness and OOD robustness.